OpenStudy (jessiegonzales):

MEDAL!!!!! 13. What is the solution to the compound inequality? -8<2x-4<2

pooja195 (pooja195):

What you do to one u do to all how to we isolate the x?

pooja195 (pooja195):

Whats our first step? :)

OpenStudy (jessiegonzales):

\[-8\le2x-4<2\]

OpenStudy (jessiegonzales):

multiply?

pooja195 (pooja195):

no we need to add the 4 that would be the 1st step :) -8+4=? -4+4=? 2+4=?

OpenStudy (jessiegonzales):

-4 0 6

pooja195 (pooja195):

\[\huge~\rm~-4<2x<6\]

OpenStudy (jessiegonzales):

oh ieee

pooja195 (pooja195):

one last step what is it? we need to isolate x

OpenStudy (jessiegonzales):

get x alone

pooja195 (pooja195):

yea how would u do that?

OpenStudy (jessiegonzales):

subtract?

pooja195 (pooja195):

We need to divide it

pooja195 (pooja195):

\[\huge~\rm~\frac{ 4<2x<6 }{ 2~~~~~2~~~~~~2 }\]

OpenStudy (jessiegonzales):

2<x<3

OpenStudy (jessiegonzales):

-2<x<3

pooja195 (pooja195):

lol thanks for catching that mistake sorry! but yes thats correct good job buddy :) !

OpenStudy (jessiegonzales):

thanks <3

pooja195 (pooja195):

<3 :)